# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 3 Markov Chain Fundamentals
Consider a function $f:\Omega\rightarrow\mathbb{R}$. The expectation of $f$ with respect to the stationary distribution $\mu$ is the following:
@@
Moreover, the variance of $f$ with respect to $\mu$ has the following equivalent formulations that can be easily verified by simple algebra:
$\mathrm{Var}_{\mu}(f):=\mathbb{E}_{\mu}[(f-\mu(f))^{2}]=\sum_{\sigma\in\Omega}\mu(\sigma)\left(f(\sigma)-\mu(f)\right)^{2}=\frac{1}{2}\sum_{\sigma,\eta\in\Omega}\mu(\sigma)\mu(\eta)(f(\sigma)-f(\eta))^{2},$
where the last sum is over all ordered pairs (see [Jer03, (5.4-5.5)] for a proof of the last identity).
Notice that the later formulation of variance is a sum over all pairs of states of the difference in their functional value (squared to measure the absolute value).
In contrast to the variance which measures the “global variation” of the function $f$, the Dirichlet form is a measure of the “local variation” for a specified Markov chain. The terms global vs. local are with respect to the graph $(\Omega,P)$.
The Dirichlet form measures the difference in the functional value over pairs of adjacent states with respect to the graph $(\Omega,P)$ as formulated in the following definition.
@@
For an ergodic Markov chain $(\Omega,P,\pi)$, recall that the stationary distribution is an eigenvector with eigenvalue $1$. Hence, we denote the eigenvalues in the following manner:
$11>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-$ 1 .
The fact that $\lambda_{N}>-1$ follows from the aperiodicity of an ergodic Markov chain. The spectral gap is defined as $\gamma=1-\lambda_{2}$. Moreover, the absolute spectral gap is $1-\max\{\lambda_{2},|\lambda_{N}|\}$. When the chain has self-loop probabilities $\geq 1/2$ (i.e., $P(\sigma,\sigma)\geq 1/2$ for all $\sigma\in\Omega$), then all eigenvalues are non-negative and hence the spectral gap equals the absolute spectral gap. The absolute spectral gap measures the decay rate of variance whereas the spectral gap is potentially off by a factor of $1/2$. However the spectral gap is more convenient to work with due to the following formulation known as a Poincare inequality (see [LP17, Lemma 13.12] for a proof of the equivalence).
@@
The spectral gap $\gamma$ is the largest constant, so that the following inequality holds for any function $f:\Omega\rightarrow\mathbb{R}$,
$\gamma\cdot\mathrm{Var}(f)\leq\mathcal{E}_{P}(f)\,.$ (2)
The spectral gap measures the decay rate of the variance of the lazy version of $P$ which is defined as:
$P_{zz}=\frac{1}{2}(P+I),$
where $I$ is the identity matrix which has 1 on the main diagonal and 0 on the off-diagonal entries. This corresponds to adding a self-loop of $1/2$ to every state $i\in\Omega$, i.e., $P_{zz}(i,i)\geq 1/2$. Notice that the eigenvalues of $P_{zz}$ lie between 00 and $1$ and the absolute spectral gap is the same as the spectral gap in this case. Hence, we have the following (see Mihail [Mih89, Jer03]):
$\mathrm{Var}(P_{zz}f)\leq(1-\gamma/2)\mathrm{Var}(f).$
This yields the following bound on the relaxation time and the mixing time [Jer95, Cap23]:
$T_{\text{relax}}(P_{zz})\leq{2n}/{\gamma}\ \ \ \mbox{ and }T_{\text{mix}}(P_{zz})\leq\frac{2n}{\gamma}\log(1/\mu^{*}),$ (3)
where $\mu^{*}=\min_{x\in\Omega}\mu(x)$.
If the absolute spectral gap is the same as the spectral gap (which is true, for example, when $P$ is positive semidefinite) then:
$T_{\text{relax}}(P)\leq{n}/{\gamma}\ \ \ \mbox{ and }T_{\text{mix}}(P)\leq\frac{n}{\gamma}\log(1/\mu^{*}).$ (4)
Moreover, Dyer, Greenhill, and Ullrich [DGU14] proved that the heat-bath block dynamics is positive semidefinite (that is, all the eigenvalues of $P$ are non-negative) and the Glauber dynamics is a special case of the heat-bath block dynamics where the blocks are single vertices. Hence, we can apply the bounds in Eq. 4 to the Glauber dynamics, in particular to obtain Theorem 2.

### v4

@@ 3.3 Dirichlet Form and Variance
Consider a function $f:\Omega\rightarrow\mathbb{R}$. The expectation of $f$ with respect to the stationary distribution $\mu$ is denoted as follows:
@@
We will also study the variance of $f$ with respect to $\mu$ which has the following convenient, equivalent formulations that can be easily verified by simple algebra:
$\mathrm{Var}_{\mu}(f)$ $:=\mu\left(\left(f-\mu(f)\right)^{2}\right)$
$=\sum_{\sigma\in\Omega}\mu(\sigma)\left(f(\sigma)-\mu(f)\right)^{2}$
$=\frac{1}{2}\sum_{\sigma,\eta\in\Omega}\mu(\sigma)\mu(\eta)(f(\sigma)-f(\eta))^{2},$
see [Jer03, (5.4-5.5)] for a proof of the last identity. Notice that the last formulation of variance is a sum over all pairs of states of the difference in their functional value (squared to measure the absolute value).
It will be convenient later to use inner product and $2$-norm to represent the variance functional. For two functions $f,g:\Omega\rightarrow\mathbb{R}$, the inner product of $f$ and $g$ with respect to $\mu$ is defined as
$\left\langle f,g\right\rangle_{\mu}:=\sum_{\eta\in\Omega}\mu(\eta)f(\eta)g(\eta)=\mu(fg),$
and the $2$-norm of $f$ with respect to $\mu$ is defined as
$\left\lVert f\right\rVert_{2,\mu}=\sqrt{\left\langle f,f\right\rangle_{\mu}}.$
Thus, we can represent the variance of $f$ as
$\mathrm{Var}_{\mu}(f)=\left\lVert f-\mu(f)\right\rVert_{2,\mu}^{2}.$
The variance functional is closely related to the $\chi^{2}$-divergence. Specifically, for any distribution $\nu$ over $\Omega$, the $\chi^{2}$-divergence of $\nu$ from $\mu$ is defined as
$D_{\chi^{2}}(\nu\,\|\,\mu):=\sum_{\eta\in\Omega}\mu(\eta)\left(\frac{\nu(\eta)}{\mu(\eta)}-1\right)^{2}=\sum_{\eta\in\Omega}\frac{\left(\nu(\eta)-\mu(\eta)\right)^{2}}{\mu(\eta)}.$
Observe that $D_{\chi^{2}}(\nu\,\|\,\mu)\geq 0$, with an equality iff $\nu=\mu$. If we let $f=\nu/\mu$ be the relative density of $\nu$ with respect to $\mu$, i.e., $f(\eta)=\nu(\eta)/\mu(\eta)$ for each $\eta\in\Omega$, then it holds $\mu(f)=1$ and $\mathrm{Var}_{\mu}(f)=D_{\chi^{2}}(\nu\,\|\,\mu)$.
In contrast to the variance which measures the “global variation” of the function $f$, the Dirichlet form is a measure of the “local variation” for a specified Markov chain. The terms global vs. local are with respect to the graph $(\Omega,P)$. The Dirichlet form measures the difference in the functional value over pairs of adjacent states with respect to the graph $(\Omega,P)$ as formulated in the following definition.
@@
For a Markov chain on state space $\Omega$ with transition matrix $P$, and for a function $f:\Omega\rightarrow\mathbb{R}$, define the Dirichlet form as follows:
@@
Equivalently, we can write the Dirichlet form as
$\mathcal{E}_{P}(f)=\left\langle f,(I-P)f\right\rangle_{\mu},$
where $I$ is the identity matrix, and (IP)f(I-P)f is a function from $\Omega$ to $\mathbb{R}$ obtained as a matrix-vector product of IPI-P and $f$ (viewed as a column vector).
3.4 Spectral Gap
For an ergodic, reversible Markov chain $(\Omega,P,\mu)$, recall that $P$ denotes the transition matrix and $\mu$ is the stationary distribution which means that $\mu$ is a row eigenvector of $P$ with eigenvalue $1$. Since the Markov chain is reversible then the eigenvalues of $P$ are real-valued, and hence we denote the eigenvalues of $P$ as follows:
$1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{N}>-1.$
The fact that $\lambda_{N}>-1$ follows from the aperiodicity of an ergodic Markov chain. The spectral gap is defined as $\gamma=1-\lambda_{2}$, and the absolute spectral gap is $1-\lambda_{*}$ where $\lambda_{*}=\max\{\lambda_{2},|\lambda_{N}|\}$ is the second largest eigenvalue in absolute value.
Dyer, Greenhill, and Ullrich [DGU14] proved that the heat-bath block dynamics is positive semidefinite (PSD), that is, all of the eigenvalues of $P$ are non-negative. The heat-bath block dynamics is defined by a collection of blocks $B_{1},\dots,B_{L}$ where each $B_{i}\subset V$ and $\cup_{i}B_{i}=V$; the steps $X_{t}\rightarrow X_{t+1}$ of the block dynamics operate by first choosing a random $B_{i}$ and then choosing the configuration $X_{t+1}(B_{i})$ on $B_{i}$ from the Gibbs distribution conditional on $X_{t}(V\setminus B_{i})$ (the configuration on $V\setminus B_{i}$ stays the same). The transition matrix of heat-bath update for one block is positive semidefinite (since it is a block diagonal matrix with positive semidefinite blocks of rank 1). The positive semidefinitness now follows since the heat bath chain is a mixture of block updates. The Glauber dynamics is the special case of the heat-bath block dynamics where the blocks are single vertices. Hence, for the Glauber dynamics, the absolute spectral gap is the same as the spectral gap.
The absolute spectral gap measures the decay rate of variance whereas the spectral gap is potentially off by a factor of $1/2$ (though this distinction is irrelevant for the case of the Glauber dynamics). The spectral gap is convenient to work with due to the following formulation known as a Poincaré inequality (see [LP17, Lemma 13.7] for a proof of the equivalence).
@@ 3.4 Spectral Gap
For a reversible Markov chain defined by $(\Omega,P,\mu)$, the spectral gap $\gamma$ is the largest constant such that the following inequality holds for any function $f:\Omega\rightarrow\mathbb{R}$,
$\gamma\cdot\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P}(f)\,.$ (15)
For non-reversible Markov chains, we can define $\gamma$ as in Eq. 15 which is referred to as the Poincaré constant (see [Jer95, Remark 5.3]); however it is not necessarily equivalent to the spectral gap $1-\lambda_{2}$ in the case of non-reversible chains.
Recall, the relaxation time is defined as the inverse of the absolute spectral gap:
$T_{\mathrm{relax}}:=(1-\lambda_{*})^{-1}.$
Let $P_{\textsc{gd}}$ denote the transition matrix of the Glauber dynamics. As noted above, for the Glauber dynamics the transition matrix $P_{\textsc{gd}}$ is PSD and hence the absolute spectral gap is the same as the spectral gap, and thus we have:
$T_{\mathrm{relax}}(P_{\textsc{gd}})\leq\gamma^{-1}.$ (16)
3.5 Approximate Tensorization of Variance
A useful tool for bounding the spectral gap of the Glauber dynamics is the following property of the Gibbs distribution known as approximate tensorization of variance.
Definition 3.4.
For a graph $G=(V,E)$, let $\mu$ be a probability distribution on $\Omega\subseteq\{0,1\}^{V}$. For a constant $C\geq 1$, we say that $\mu$ satisfies $C$-approximate tensorization of variance if the following holds for all $f:\Omega\rightarrow\mathbb{R}$,
$\mathrm{Var}_{\mu}(f)\leq C\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)].$ (17)
It is important to stress that approximate tensorization of variance is a property of the Gibbs distribution and does not explicitly involve the Glauber dynamics, but as we will see shortly it is equivalent to bounding the spectral gap. When $\mu$ is a product distribution, i.e., $\mu=\otimes_{v\in V}\mu(v)$, then Eq. 17 holds with $C=1$, which is the optimal constant (it is necessary that $C\geq 1$, as can be seen by taking functions depending only on a single vertex).
Let us decipher the right-hand side of Eq. 17 before proceeding: for the term $\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]$, we first sample a configuration $\tau$ on $V\setminus\{v\}$ where $\tau$ is drawn from the projection of $\mu$ on $V\setminus\{v\}$, then we consider the variance of the function $f$ with respect to the marginal distribution of $\mu$ on $v$ conditional on $\tau$. More formally, for a configuration $\tau\in\{0,1\}^{V\setminus\{v\}}$ on all vertices except $v$, let $\mu(\tau)$ denote the marginal probability of $\tau$ in the Gibbs distribution $\mu$, i.e., $\mu(\tau)=\sum_{\sigma\in\Omega:\sigma(V\setminus\{v\})=\tau}\mu(\sigma)$. Conversely, let $\mu_{\tau}$ denote the Gibbs distribution $\mu$ conditional on $\tau$, and thus for $i\in\{0,1\}$,
$\mu_{\tau}(\eta(v)=i)=\mu(\eta(v)=i\mid\eta(V\setminus\{v\})=\tau).$
Then we have the following:
$\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]=\sum_{\tau\in\{0,1\}^{V\setminus\{v\}}}\mu(\tau)\mu_{\tau}(\eta(v)=0)\mu_{\tau}(\eta(v)=1)\left(f(\tau_{v,0})-f(\tau_{v,1})\right)^{2},$ (18)
where, for $i\in\{0,1\}$, $\tau_{v,i}$ is the extension of $\tau$ to an assignment on all vertices where for $w\in V$:
$\tau_{v,i}(w):=\begin{cases}i,&\mbox{if }w=v;\\ \tau(w),&\mbox{if }w\neq v.\end{cases}$
For a configuration $\sigma\in\Omega$ on all vertices $V$, for $v\in V$, let $\sigma_{v}=\sigma_{v,1-\sigma(v)}$ denote the configuration which flips the spin at $v$ and keeps all other spins the same. Let $P_{\textsc{gd}}$ denote the transition matrix for the Glauber dynamics. Using the observation that $\mu^{\sigma(V\setminus\{v\})}(\eta(v)=1-\sigma(v))=nP_{\textsc{gd}}(\sigma,\sigma_{v})$ and recalling the definition of the Dirichlet form (see Definition 3.2), we can rewrite Eq. 18 as follows:
$\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]$ $=\frac{n}{2}\sum_{v\in V}\sum_{\sigma\in\Omega}\mu(\sigma)P_{\textsc{gd}}(\sigma,\sigma_{v})\left(f(\sigma)-f(\sigma_{v})\right)^{2}$ (19)
$=n\mathcal{E}_{P_{\textsc{gd}}}(f).$ (20)
Therefore,
$\mathrm{Var}_{\mu}(f)\leq C\sum_{v\in V}\mathbb{E}_{\mu}[\mathrm{Var}_{v}(f)]\quad\iff\quad\frac{1}{Cn}\cdot\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P}(f),$
and hence $C$-approximate tensorization of variance is equivalent to the spectral gap satisfying $\gamma\geq 1/(Cn)$.
Corollary 3.5.
A distribution $\mu\in\{0,1\}^{V}$ satisfies $C$-approximate tensorization of variance iff the Glauber dynamics for $\mu$ has spectral gap $\gamma\geq 1/(Cn)$ (or equivalently, the Glauber dynamics has relaxation time $T_{\mathrm{relax}}\leq Cn$).
The analog of approximate tensorization of variance with entropy in place of variance is presented in Section 6.8, see Definition 6.8. Whereas approximate tensorization of variance implies optimal bounds of $O(n)$ time on the relaxation time, the stronger condition of approximate tensorization of entropy implies optimal bounds of $O(n\log{n})$ time on the mixing time. We refer the interested reader to Section 6.8 for further details.
3.6 Preliminaries
Here we present a functional characterization of spectral independence, utilizing the variance functional.
Definition 3.6.
We say that a measure $\mu$ on $\{0,1\}^{n}$ satisfies approximate subadditivity of variance with constant $C$ if for every $f:\Omega\to\mathbb{R}$ we have
$\sum_{i=1}^{n}\mathrm{Var}_{\mu}[\mathbb{E}[f|X_{i}]]\leq C\cdot\mathrm{Var}_{\mu}[f].$ (21)
Lemma 3.7.
Measure $\mu$ satisfies $\lambda_{\max}(\Psi_{\mu})\leq C$ if and only if $\mu$ satisfies approximate subadditivity of variance with constant $C$.
Proof.
Without loss of generality, we can assume $\mathbb{E}_{\mu}[f]=0$. Let $p_{i}=\mathrm{Pr}(X_{i}=1)\in(0,1)$ for $i\in[n]$. Let $a_{i}=\mathbb{E}[f|X_{i}=1]$ and note that $\mathbb{E}[f|X_{i}=0]=a_{i}p_{i}/(1-p_{i})$, since $\mathbb{E}_{\mu}[f]=0$. In terms of $p_{i}$ and $a_{i}$’s the LHS of (21) is
$\sum_{i=1}^{n}a_{i}^{2}\frac{p_{i}}{1-p_{i}}.$ (22)
We only need to verify (21) for $f$ that minimizes $\mathrm{Var}_{\mu}[f]$ subject to $\mathbb{E}[f|X_{i}=1]=a_{i}$ for $i\in[n]$ and $\mathbb{E}_{\mu}[f]=0$. From the method of Lagrange multipliers it follows that such $f$ is linear, that is, for some $c_{0},c_{1},\dots,c_{n}$ we have
$f=c_{0}+\sum_{i=1}^{n}c_{i}X_{i}.$
Let $c=(c_{1},\dots,c_{n})$. In terms of $c_{i}$’s the main term on the RHS of (21) is
$\mathrm{Var}_{\mu}[f]=c^{T}\mathrm{Cov}_{\mu}c.$
The $c_{i}$’s, $a_{i}$’s, and $p_{i}$’s satisfy the following equation (for $i\in[n]$):
$(\mathrm{Cov}_{\mu}c)_{i}=\mathbb{E}[fX_{i}]-\mathbb{E}[f]\mathbb{E}[X_{i}]=\mathrm{Pr}[X_{i}=1]\mathbb{E}[f|X_{i}=1]-0=p_{i}a_{i}.$ (23)
Let $D=\mathrm{diag}(\mathrm{Var}(X_{i}))$ be the variance matrix. We have $D_{ii}=p_{i}(1-p_{i})$ for $i\in[n]$ and recall that $C$-spectral independence means $\mathrm{Cov}_{\mu}\preceq CD$, see Eq. 10. Thus we can express the LHS of (21) as $c^{T}\mathrm{Cov}_{\mu}D^{-1}\mathrm{Cov}_{\mu}c$ and (21) holding for every $f$ is equivalent to the following
$\mathrm{Cov}_{\mu}D^{-1}\mathrm{Cov}_{\mu}\preceq C\,\mathrm{Cov}_{\mu}.$ (24)
If $\mu$ is $C$ spectrally independent then we have $D^{-1}\preceq C\mathrm{Cov}_{\mu}^{-1}$ and by multiplying by $\mathrm{Cov}_{\mu}$ on both left and right we obtain (24) and hence $C$-approximate subadditivity of variance (technically one should use the pseudo-inverse for $\mathrm{Cov}_{\mu}$ since it is not guaranteed to be invertible; this does not change the proof).
If $\mu$ satisfies $C$-approximate subadditivity of entropy then we have (24); multiplying both sides by $\mathrm{Cov}_{\mu}^{-1}$ (technically we are using the pseudo-inverse for $\mathrm{Cov}_{\mu}$) we obtain $D^{-1}\preceq C\,\mathrm{Cov}_{\mu}^{-1}$, which implies $\mathrm{Cov}_{\mu}\preceq C\cdot D$ and hence, by Eq. 10, $\lambda_{\max}(\Psi_{\mu})\leq C$. ∎
3.7 Mixing Time Bounds via Spectral Gap
We now show how the spectral gap implies an upper bound on the mixing time. For the Glauber dynamics $(X_{t})$, let $P_{\textsc{gd}}$ denote the transition matrix, let $\nu_{t}$ denote the distribution of the random variable $X_{t}$ at time $t$, and let $\mu$ denote the stationary distribution. Let
$\mu^{*}=\min_{x\in\Omega}\mu(x).$
We will show the following bound:
$T_{\mathrm{mix}}(P_{\textsc{gd}})\leq\frac{1}{2\gamma}\ln(4/\mu^{*}).$ (25)
We will also establish a more general bound which accounts for a so-called warm start when the initial distribution $\nu_{0}$ is somewhat close to the stationary distribution $\mu$.
Let $f_{t}=\nu_{t}/\mu$ denote the relative density at time $t$, i.e., $f_{t}(x)=\nu_{t}(x)/\mu(x)$ for each $x\in\Omega$. Notice that $\mu(f_{t})=1$. Recall the $\chi^{2}$-divergence from stationarity defined as:
$D_{\chi^{2}}(\nu_{t}\,\|\,\mu)=\mathrm{Var}_{\mu}(f_{t})=\left\|f_{t}-1\right\|_{2,\mu}^{2}=\sum_{x\in\Omega}\mu(x)\left(\frac{\nu_{t}(x)}{\mu(x)}-1\right)^{2}.$
Meanwhile, the $L^{1}$-distance from stationarity is related to the total variation distance:
$\left\|f_{t}-1\right\|_{1,\mu}=\sum_{x\in\Omega}|\nu_{t}(x)-\mu(x)|=2\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}.$
We can bound the convergence time to the stationary distribution $\mu$ as a function of the initial distance. Let
$M:=D_{\chi^{2}}(\nu_{0}\,\|\,\mu)$ (26)
denote the initial $\chi^{2}$-divergence from the stationary distribution. Note that $M$ is typically exponentially large in $n$:
$M=\left(\sum_{x}\frac{\nu_{0}(x)^{2}}{\mu(x)}\right)-1\leq\frac{1}{\mu^{*}}\sum_{x}\nu_{t}(x)^{2}\leq\frac{1}{\mu^{*}}.$ (27)
In contrast, when $M=O(1)$ (e.g., if $\nu_{0}(x)\leq 2\mu(x)$ for all $x\in\Omega$) then we refer to $\nu_{0}$ as a warm start. We will prove the following bound on the distance to stationarity of the Glauber dynamics after $t$ steps as a function of the initial distance $M$:
$4\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}^{2}\leq D_{\chi^{2}}(\nu_{t}\,\|\,\mu)\leq M\exp(-2\gamma t).$ (28)
The above bound holds for the Glauber dynamics, and more generally holds for any reversible chain whose transition matrix is positive semidefinite (consequently, the absolute spectral gap is the same as the spectral gap); we will prove it below. Plugging in the worst case bound for the initial distance $M$ from Eq. 27 into Eq. 28 yields the upper bound on the mixing time stated earlier in Eq. 25.
For general Markov chains (not necessarily reversible), defined by a transition matrix $P$, recall $\gamma$ is defined in Eq. 15. Note that for non-reversible chains the eigenvalues of the transition matrix are not necessarily real. For general chains (not necessarily reversible) a similar result as in Eq. 28 holds with $\gamma/2$ in place of $2\gamma$ for the lazy version $P_{zz}:=\frac{1}{2}(P+I)$ (which transitions as in the original chain $P$ with probability $1/2$ and stays in the same state with probability $1/2$), see [Mih89] and [Jer03, Theorem 5.6].
Notice that if we have a warm start for the initial distance then the mixing time is on the order of the inverse spectral gap. Consequently, in many approximate counting algorithms, such as the simulated annealing algorithm of [ŠVV09, Hub15, Kol18, HK25], the running time avoids the extra $O(\log(1/\mu^{*}))$ factor.
The main task is to prove Eq. 28; then Eq. 25 follows as a corollary as noted above by plugging in the worst-case bound on $M$ from Eq. 27. We refer the interested reader to [Jer03, Chapter 5] for an alternative (and somewhat simpler) proof of Eq. 28 for the lazy chain $P_{zz}$. The below proof is stated for the Glauber dynamics but holds for any PSD transition matrix.
Proof of Eq. 28.
[…]
$\mathrm{Var}(P_{\textsc{gd}}f)\leq(1-\gamma)^{2}\mathrm{Var}(f),\quad\forall f:\Omega\to\mathbb{R};$
$\quad\iff\quad$ $D_{\chi^{2}}(\nu P_{\textsc{gd}}\,\|\,\mu)\leq(1-\gamma)^{2}D_{\chi^{2}}(\nu\,\|\,\mu),\quad\text{$\forall$ distribution $\nu$ on $\Omega$}.$ (29)
We will establish Eq. 29 later in the proof. Note, for the analog of Eq. 29 for the lazy version of general chains (including non-reversible chains), see [Mih89] and, e.g., [Jer03, Corollary 5.7].
To complete the proof of Eq. 28, assuming Eq. 29 holds, we will begin by showing the first inequality in Eq. 28. Namely, by Cauchy–Schwarz inequality the $L^{1}$-distance is upper bounded by the $L^{2}$-distance (see also [MT06, (7.1)]):
$2\left\|\nu_{t}-\mu\right\|_{\mathrm{TV}}=\left\|f_{t}-1\right\|_{1,\mu}=\mu\left(|f_{t}-1|\right)\leq\sqrt{\mu\left((f_{t}-1)^{2}\right)}=\left\|f_{t}-1\right\|_{2,\mu}=\sqrt{D_{\chi^{2}}(\nu_{t}\,\|\,\mu)}.$
For the second inequality in Eq. 28, we deduce from Eq. 29 that
$D_{\chi^{2}}(\nu_{t}\,\|\,\mu)=D_{\chi^{2}}(\nu_{0}P^{t}\,\|\,\mu)\leq(1-\gamma)^{2t}D_{\chi^{2}}(\nu_{0}\,\|\,\mu)\leq M\exp(-2\gamma t),$
where $M$ is the $\chi^{2}$-divergence of the initial distribution $\nu_{0}$ to the stationary distribution $\mu$, see Eq. 26.
It remains to prove Eq. 29. We claim the following holds:
$\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\mathrm{Var}_{\mu}(f)-\mathcal{E}_{P_{\textsc{gd}}^{2}}(f).$ (30)
Without loss of generality, we assume $\mu(f)=0$. We observe that
$\mathrm{Var}_{\mu}(f)=\left\lVert f\right\rVert_{2,\mu}^{2}=\langle f,f\rangle_{\mu}$
and
$\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\left\lVert P_{\textsc{gd}}f\right\rVert_{2,\mu}^{2}=\langle P_{\textsc{gd}}f,P_{\textsc{gd}}f\rangle_{\mu}=\langle f,P_{\textsc{gd}}^{2}f\rangle_{\mu},$
where the last equality follows from the reversibility of $P_{\textsc{gd}}$. Therefore, we deduce that
$\mathrm{Var}_{\mu}(f)-\mathrm{Var}_{\mu}(P_{\textsc{gd}}f)=\langle f,(I-P_{\textsc{gd}}^{2})f\rangle_{\mu}=\mathcal{E}_{P_{\textsc{gd}}^{2}}(f).$
By the Poincaré inequality Eq. 15 $\gamma\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P_{\textsc{gd}}}(f)$, we have that $(1-(1-\gamma)^{2})\mathrm{Var}_{\mu}(f)\leq\mathcal{E}_{P_{\textsc{gd}}^{2}}(f)$ since every eigenvalue of $P_{\textsc{gd}}^{2}$ is the square of an eigenvalue of $P_{\textsc{gd}}$, and the Glauber dynamics $P_{\textsc{gd}}$ is positive semidefinite. From Eq. 30, we then obtain Eq. 29 that
$\mathrm{Var}(P_{\textsc{gd}}f)\leq(1-\gamma)^{2}\mathrm{Var}(f),$
completing the proof. ∎
We showed in Eq. 25 how a lower bound on the spectral gap of the Glauber dynamics (which is equivalent to an upper bound on the relaxation time) implies an upper bound on the mixing time with an extra $O(\log(1/\mu^{*}))$ factor. In Remark 9.30 in Section 9.5 we discuss the reverse connection: namely, an optimal upper bound on the mixing time of the Glauber dynamics of the form $T_{\mathrm{mix}}(\varepsilon)=O(n\log(n/\varepsilon))$ for all $\varepsilon>0$ implies an optimal upper bound on the relaxation time of $T_{\mathrm{relax}}=O(n)$.